Calculate the number of atoms of your element that has a mass equal to your birth date + your birth month (ex. My birthday is Dec. 19; 12 + 19 = 31g). Show all of your work and express your final answer with the correct number of significant figures.
1
Expert's answer
2020-12-03T14:25:35-0500

Jun 6,

Total mass = 6+6 = 12g


12.00 g C-12 = 1 mol C-12 atoms = 6.022 × 1023 atoms • The number of particles in 1 mole is called Avogadro's Number (6.0221421 x 1023).